# FlakyFix: Using Large Language Models for Predicting Flaky Test Fix Categories and Test Code Repair

### v3

@@ 5.2 Internal Threats
Internal threats to validity are about to make sure the cause-effect relation identified in the study is really there and there is no other explanation (confounding factors). In our study, the claim is that  our predicted fix categories will help fix flakiness. The main confounding factor is about the way we prompt the GPT-3.5 model, when asking to fix a flaky test. Better results might be obtained in the future with additional guidance and more examples in the prompt. Similarly, using GPT-4 would probably lead to even better results but, as discussed in Section 2.3, this would raise, on our data set, possible data leakage issues.
@@ 7 Conclusion and Future Work
In this paper, we proposed a framework to (a) categorize flaky tests according to practical fix categories based exclusively on test code and (b) use the fix categories to generate fully automated repairs of the flaky tests with language models. Given the limited availability of open-source datasets for flaky test research, our generated labeled data can also be valuable for future research, particularly for ML-based prediction and repair of flaky tests. We evaluated two language models, CodeBERT and UniXcoder, with and without Few-Shot Learning, and found that the UniXcoder model outperformed CodeBERT in predicting most fix categories. We also showed that the category information, complemented with in-context learning can greatly help a large language model (LLM), namely GPT-3.5 in our case, better fix the flakiness in test code. Lastly, we executed a sample of GPT-repaired flaky tests and estimated that a large percentage (roughly between 70% and 90%) of these tests can be expected to pass. For the generated tests that fail, an average of 16% of the code needs further modification for them to pass. In the future, we plan to expand our flaky dataset with more data. We also plan to expand our heuristics for automatically labeling flaky tests to other programming languages. Replicating the study with other LLMs (such as GPT-4.0) will also be made possible with an updated dataset. Additionally, a significant percentage of flaky tests may stem from problems in the production code, which cannot be addressed by black-box models. Therefore, in the future, we need to develop lightweight and scalable approaches for accessing and fixing the production code when dealing with test flakiness.

### v4

@@ 5.2 Internal Threats
Internal threats to validity are about to make sure the cause-effect relation identified in the study is really there and there is no other explanation (confounding factors). In our study, the claim is that  our predicted fix categories will help fix flakiness. The main confounding factor is about the way we prompt the GPT-3.5 model, when asking to fix a flaky test. Better results might be obtained in the future with additional guidance and more examples in the prompt. Similarly, using GPT-4 would probably lead to even better results but, as discussed in Section 2.3, this would raise, on our data set, possible data leakage issues. Additionally, it would be insightful to analyze how often GPT generates a correct fix when provided with an incorrectly predicted fix category. However, as explained in Section 3.2.5, we were limited to executing only 35 tests. Within our sample of 35 executable test cases, we encountered only one instance where the fix category was incorrectly predicted, albeit partially correct as one of the two categories was accurately predicted. Therefore, conducting this analysis reliably is not possible with our current data.
@@ 7 Conclusion and Future Work
In this paper, we proposed a framework to (a) categorize flaky tests according to practical fix categories based exclusively on test code and (b) use the fix categories to generate fully automated repairs of the flaky tests with language models. Given the limited availability of open-source datasets for flaky test research, our generated labeled data can also be valuable for future research, particularly for ML-based prediction and repair of flaky tests. We evaluated two language models, CodeBERT and UniXcoder, with and without Few-Shot Learning, and found that the UniXcoder model outperformed CodeBERT in predicting most fix categories. We also showed that the category information, complemented with in-context learning can greatly help a large language model (LLM), namely GPT-3.5 in our case, better fix the flakiness in test code. Lastly, we executed a sample of GPT-repaired flaky tests and estimated that a large percentage (roughly between 51% and 83%) of these tests can be expected to pass. For the generated tests that fail, an average of 16% of the code needs further modification for them to pass. In the future, we plan to expand our flaky dataset with more data. We also plan to expand our heuristics for automatically labeling flaky tests to other programming languages. Replicating the study with other LLMs (such as GPT-4.0) will also be made possible with an updated dataset. Additionally, a significant percentage of flaky tests may stem from problems in the production code, which cannot be addressed by black-box models. Therefore, in the future, we need to develop lightweight and scalable approaches for accessing and fixing the production code when dealing with test flakiness.